# Stabilizing and Solving Inverse Problems using Data and Machine Learning

### v1

@@ Abstract
We consider an inverse problem involving the reconstruction of the solution to a nonlinear partial differential equation (PDE) with unknown boundary conditions. Instead of direct boundary data, we are provided with a large dataset of boundary observations for typical solutions (collective data) and a bulk measurement of a specific realization. To leverage this collective data, we first compress the boundary data using proper orthogonal decomposition (POD) in a linear expansion. Next, we identify a possible nonlinear low-dimensional structure in the expansion coefficients using an auto-encoder, which provides a parametrization of the dataset in a lower-dimensional latent space. We then train a neural network to map the latent variables representing the boundary data to the solution of the PDE. Finally, we solve the inverse problem by optimizing a data-fitting term over the latent space.
We analyze the underlying stabilized finite element method in the linear setting and establish optimal error estimates in the $H^{1}$ and $L^{2}$-norms. The nonlinear problem is then studied numerically, demonstrating the effectiveness of our approach.
@@ 1 Introduction
This is the situation that we wish to address in the present work. The objective is to construct a computational method that combines machine learning techniques for the data handling parts and hybrid network/finite element methods for the approximation in physical space. First, the large data set is mapped to a lower dimensional manifold using an auto-encoder or some other technique for finding low dimensional structures such as singular value decomposition or manifold learning. Then, we train a network to reproduce the solution map from the lower dimensional set to the finite element space. Finally, this reduced order model solves a non-linear inverse problem under the a priori assumption that the solution resides in a neighborhood of the lower dimensional manifold.
To ensure an underpinning of the developed methods, we consider the case of a unique continuation problem for a non-linear elliptic operator. That is, given some interior measurement (or measurements on the part of the boundary), a solution is reconstructed despite lacking boundary data on the part of the boundary. Such problems are notoriously ill-posed, and using only the event data set, it is known that the accuracy of any approximation in the whole domain cannot be guaranteed due to the poor global stability [1]. Indeed, in general, stability is no better than logarithmic. This means that for perturbations of order $\epsilon$, the error must be expected to be of order $|\log(\epsilon)|^{-\alpha}$ with $\alpha\in(0,1)$. In interior subdomains stability is of Hölder type, meaning that the same perturbation gives rise to an $O(\epsilon^{\alpha})$ error. Computational methods can have, at best, rates that reflect this stability of the continuous problem [8]. A convenient a priori assumption is that the missing data of the approximate solution is in a $\delta$-neighbourhood of a finite-dimensional space, $D_{N}$, where $\delta$ is the smallest distance from the solution to $D_{N}$ in some suitable topology. In this case, it is known that the stability is Lipschitz; that is, the problem has similar stability properties to a well-posed problem, and finite element methods can be designed with optimal convergence up to the data approximation error $\delta$. For linear model problems discretized using piecewise affine finite element methods with mesh parameter $h$, one can prove the error bound [9],
@@
Here, $C_{N}$ is a constant that depends on the dimension of the data set $D_{N}$, the geometry of the available event data, and the smoothness of the exact solution. In particular, $C_{N}$ typically grows exponentially in $N$.
Since the size of $N$ must be kept down, there is a clear disadvantage in using the full large dataset. Indeed, for $N$ sufficiently large, the experience data will have no effect. Instead, we wish to identify a lower-dimensional structure in the high-dimensional dataset, a lower-dimensional manifold such that the data resides in a $\delta$-neighbourhood of the manifold. For this task, one may use proper orthogonal decomposition in the linear case or neural network auto-encoders in the general case.
@@
  • We prove that the inverse problem with boundary data in a finite-dimensional set $\mathcal{G}$ is stable and design a method that reconstructs the solution using the reduced order basis with the same dimension as $\mathcal{G}$. We proved optimal error bounds in the $H^{1}$-norm for this method, where the constant of the error bound grows exponentially with the dimension of $\mathcal{G}$.
  • In the situation where a large set of perturbed random data, $\mathcal{G}_{S}$, from the set $\mathcal{G}$ is available, we develop a practical method for the solution of the severely ill-posed inverse problem of unique continuation, leveraging the large dataset to improve the stability properties. In order to handle non-linearity in the PDE-operator and data efficiently we adopt machine learning algorithms. The machine learning techniques are used for the following two subproblems:
@@
      Identification of the latent space $\mathcal{G}$ from $\mathcal{G}_{S}$ to find the smallest possible space for the inverse identification.
@@
      Construction of the solution map
      $\phi_{u,N,h}:G_{N}\rightarrow V_{h}\subset H^{1}(\Omega)$ (1.1)
      which approximates the Galerkin approximation of
      $\mathcal{P}(u)=0\quad\text{in $\Omega$},\qquad u=\phi_{F,N}^{-1}(\widehat{g})\quad\text{on $\partial\Omega$}$ (1.2)
      where $\mathcal{P}(u)$ is the non-linear PDE in question.
@@
The inverse problem we consider herein is of unique continuation type. There are many types of methods for this type of problem. In the framework we consider the earliest works considered quasi-reversibility [3]. The stabilized method we consider for unique continuation was first proposed in [4, 5] and [6]. More recent works use residual minimization in dual norm [11, 13, 7]. The optimal error estimates for unique continuation with a trace in a finite dimensional space was first considered for Dirichlet trace in [9] and for Neumann trace in [10]. The idea of combining unique continuation in finite dimensional space with collective data was first proposed in [19, 2] using linear algebra methods for the compression and direct solution of the linear unique continuation problem. Low rank solvers for the solution of inverse problems have also been designed in [26] using proper orthogonal decomposition.
In recent years, significant advancements have been made in utilizing machine learning for solving PDEs and efficiently representing the solutions or deriving reduced order models [16, 20, 29, 23, 27, 30]. These developments are very useful in the context of inverse problems, where they have been utilized in both data- and model-driven inverse problems [24, 25, 21, 18, 17, 15, 12].
@@
In Section 2, we introduce the model problem and the finite element discretization; in Section 3, we present and prove stability and error estimates; in Section 4, we develop a machine learning-based approach for solving the inverse problem; in Section 5, we present several numerical examples illustrating the performance of the method for various complexity of the given set of boundary data; and in Section 6 we summarize our findings and discuss future research directions.
@@ 2.1 Inverse Problem
$\boxed{\inf_{v\in V}\frac{1}{2}\|u_{0}-v\|_{H^{1}(\omega)}^{2}\quad\text{subject to}\quad\mathcal{P}(v)=0\quad\text{in $\Omega$}}$ (2.1)
@@
of observed Dirichlet data at the boundary. The dataset $\mathcal{G}$ may have different properties, but here we will assume that is of the form
@@
$\boxed{\inf_{v\in V}\frac{1}{2}\|u_{0}-v\|_{H^{1}(\omega)}^{2}\quad\text{subject to}\quad\mathcal{P}(v)=0\quad\text{in $\Omega$},\quad v|_{\partial\Omega}\in\mathcal{G}}$ (2.5)
@@ 2.2 Finite Element Method
for $10\leq m\leq k\leq p+$ 1. Here $N(T)$ is the union of all elements that share a node with $T$.
@@
$\boxed{\inf_{v\in V_{h}}\frac{1}{2}\|u_{0}-v\|_{H^{1}(\omega)}^{2}\quad\text{subject to}\quad(\mathcal{P}(v),w)_{H^{-1}(\Omega),H^{1}(\Omega)}=0\quad\forall v\in V_{h,0},\quad v|_{\partial\Omega}\in\pi_{h}\mathcal{G}}$ (2.7)
@@ 3.1 The Continuous Problem
$g=\mathcal{G}_{N}=\Bigl{\{}\sum_{n=1}^{N}a_{n}g_{n}\,|\,a_{n}\in\mathbb{R}\Bigr{\}}$ (3.2)
@@
$\mathcal{P}\varphi_{n}=0\quad\text{in $\Omega$},\qquad\varphi_{n}=g_{n}\quad\text{on $\Omega$}$ (3.3)
@@
By unique continuation for harmonic functions, we conclude that $u_{1}-u_{2}$ is zero on the boundary and therefore $u_{1}=u_{2}$ since the set $\{g_{n}\}_{n=1}^{N}$ is linearly independent on $\partial\Omega$. It follows that $\{\varphi_{n}\}_{n=1}^{N}$ is linearly independent on $\omega$ and by finite dimensionality, there is a constant such that
@@ 3.2 The Discrete Problem
The estimate (3.15) follows directly using the Cauchy-Schwarz inequality and the error estimates (3.12) for the approximate basis
@@ 3.3 Error Estimates
Let $m_{h}$ be defined by (3.23). Then, there is a constant such that
@@
    where we finally used the identity $I_{h}\widehat{v}=v$, which holds since $v\in\mathcal{U}_{N,h}$. Next, we bound the the second term using the stabilizing terms in $m_{h}$. To that end, we observe that we have the orthogonality
@@
    we obtain
@@
    $\|e\|_{\omega}\lesssim h|u|_{H^{2}(\Omega)}$ (3.56)
@@
Observe that the stabilization is never explicitly used in order to obtain error estimates. Indeed its only role is to ensure the bound $\|\widehat{u}_{N,h}\|_{\mathbb{R}^{N}}\lesssim\|u_{N,h}\|_{m_{h}}$ without condition on the mesh. On the other hand, due to the stabilization, it appears hard to obtain $L^{2}$-error estimates. If indeed the mesh is so fine that the optimization problem is well posed without stabilization, then we obtain error bounds in a similar fashion, both for the $H^{1}$-norm and the $L^{2}$-norm.
Example.
@@ 4 Methods Based on Machine Learning
  • Construct a parametrization of the data set by first approximately expanding the samples in a finite series of functions, for instance, using Proper Orthogonal Decomposition, and secondly using an autoencoder to find a possible nonlinear low dimensional structure in the coefficients.
@@
  • Composing the decoder, which maps the latent space to expansion coefficients with the solution network, we obtain a differentiable mapping that can be used to solve the inverse problem efficiently.
@@ 4.1 Processing the Boundary Data
where $\widehat{g}_{n}=(g,\widehat{g}_{n})_{\mathbb{R}^{N}}$. We introduce the mapping
@@
We have
@@
Next, we seek to find a possible nonlinear low dimensional structure in the POD coefficients using an autoencoder
@@
trained to minimize the loss
$\frac{1}{|G_{N}|}\sum_{\widehat{g}\in G_{N}}\|\widehat{g}-\phi_{{\text{de}}}\circ\phi_{{\text{en}}}\widehat{g}\|^{2}_{\mathbb{R}^{N}}$ (4.6)
See Figure 2(a) for a schematic illustration. Here $Z\sim\mathbb{R}^{n_{Z}}$ is the latent space with dimension $n_{Z}<N$. If there is a low dimensional structure, we may often take $n_{Z}$ significantly lower than $N$.
@@ 4.2 Operator Learning
Next, we discretize the problem using finite elements and train a network
@@
which approximates the Galerkin approximation of
$\mathcal{P}(u)=0\quad\text{in $\Omega$},\qquad u=\phi_{F,N}^{-1}(\widehat{g})\quad\text{on $\partial\Omega$}$ (4.8)
see Figure 2(b). The output of the network is the finite element degrees of freedom (DoFs). For the training of the network we use the energy functional $E_{\mathcal{P}}$ corresponding to the differential operator $\mathcal{P}$ as the foundation for the loss function. Letting $\mathbb{E}$ denote the expectation operator and $P$ an arbitrary probability distribution, the loss function that we minimize during training is
$\mathbb{E}_{g_{N}\sim P(g_{N})}\left[E_{\mathcal{P}}(\phi_{u,N,h}(g_{N}))\right]$ (4.9)
If there is no corresponding energy functional, one can instead minimize the residual of the finite element problem. It should be noted though, that assembling the residual instead of the energy has a greater computational cost and that the residual is not as easily and naturally decomposed into its local contributions as the energy. For technical details about network architecture and training used in this work, we refer to the example section 5.2.

### v2

@@ Abstract
We consider an inverse problem involving the reconstruction of the solution to a nonlinear partial differential equation (PDE) with unknown boundary conditions. Instead of direct boundary data, we are provided with a large dataset of boundary observations for typical solutions (collective data) and a bulk measurement of a specific realization. To leverage this collective data, we first compress the boundary data using proper orthogonal decomposition (POD) in a linear expansion. Next, we identify a possible nonlinear low-dimensional structure in the expansion coefficients using an autoencoder, which provides a parametrization of the dataset in a lower-dimensional latent space. We then train an operator network to map the expansion coefficients representing the boundary data to the finite element solution of the PDE. Finally, we connect the autoencoder’s decoder to the operator network which enables us to solve the inverse problem by optimizing a data-fitting term over the latent space. We analyze the underlying stabilized finite element method in the linear setting and establish an optimal error estimate in the $H^{1}$-norm. The nonlinear problem is then studied numerically, demonstrating the effectiveness of our approach.
@@ 1 Introduction
This is the situation that we wish to address in the present work. The objective is to construct a computational method that combines machine learning techniques for the data handling parts and hybrid network/finite element methods for the approximation in physical space. First, the large data set is mapped to a lower-dimensional manifold using an autoencoder or some other technique for finding low-dimensional structures, such as singular value decomposition or manifold learning. Then, we train a network to reproduce the solution map from the lower-dimensional set to the finite element space. Finally, this reduced order model solves a nonlinear inverse problem under the a priori assumption that the solution resides in a neighborhood of the lower-dimensional manifold.
To ensure an underpinning of the developed methods, we consider the case of a unique continuation problem for a nonlinear elliptic operator. That is, given some interior measurement (or measurements on the part of the boundary), a solution is reconstructed despite lacking boundary data on the part of the boundary. Such problems are notoriously ill-posed, and using only the event data set, it is known that the accuracy of any approximation in the whole domain cannot be guaranteed due to the poor global stability [1]. Indeed, in general, stability is no better than logarithmic. This means that for perturbations of order $\epsilon$, the error must be expected to be of order $|\log(\epsilon)|^{-\alpha}$ with $\alpha\in(0,1)$. In interior subdomains stability is of Hölder type, meaning that the same perturbation gives rise to an $O(\epsilon^{\alpha})$ error. Computational methods can have, at best, rates that reflect this stability of the continuous problem [11]. To improve on these estimates additional assumptions on the solution are needed. A convenient a priori assumption is that the missing data of the approximate solution is in a $\delta$-neighbourhood of a finite $N$-dimensional space, $\mathcal{G}$, where $\delta$ is the smallest distance from the solution to $\mathcal{G}$ in some suitable topology. In this case, it is known that the stability is Lipschitz; that is, the problem has similar stability properties to a well-posed problem, and finite element methods can be designed with optimal convergence up to the data approximation error $\delta$. For linear model problems discretized using piecewise affine finite element methods with mesh parameter $h$, one can prove the error bound [12],
@@
Here, $C_{N}$ is a constant that depends on the dimension $N$ of the data set $\mathcal{G}$, the geometry of the available event data, and the smoothness of the exact solution. In particular, $C_{N}$ typically grows exponentially in $N$.
Since the size of $N$ must be kept down, there is a clear disadvantage in using the full large dataset. Indeed, for $N$ sufficiently large, the experience data will have no effect. Instead, we wish to identify a lower-dimensional structure in the high-dimensional dataset, a lower-dimensional manifold such that the data resides in a $\delta$-neighbourhood of the manifold. For this task, one may use proper orthogonal decomposition in the linear case or neural network autoencoders in the general case.
@@
  • We prove that the inverse problem with boundary data in a finite-dimensional set $\mathcal{G}$ is stable and design a method that reconstructs the solution using the reduced order basis with the same dimension as $\mathcal{G}$. We prove optimal error bounds in the $H^{1}$-norm for this method, where the constant of the error bound grows exponentially with the dimension of $\mathcal{G}$.
  • In the situation where a large set of perturbed random data, $\mathcal{G}_{S}$, from the set $\mathcal{G}$ is available, we develop a practical method for the solution of the severely ill-posed inverse problem of unique continuation, leveraging the large dataset to improve the stability properties. In order to handle nonlinearity in the PDE operator and data efficiently we adopt machine learning algorithms. The machine learning techniques are used for the following two subproblems:
@@
      Identification of a potential latent space of $\mathcal{G}$ from $\mathcal{G}_{S}$ to find the smallest possible space for the inverse identification.
@@
      Construction of a discrete approximation of the solution map
      $\phi_{u}:\mathcal{G}\rightarrow H^{1}(\Omega)$ (1.1)
      that gives an approximation of the finite element solution to
      $\mathcal{P}(u)=0\quad\text{in $\Omega$},\qquad u|_{\partial\Omega}\in\mathcal{G}$ (1.2)
      where $\mathcal{P}$ is the nonlinear PDE operator in question. The construction is done in a way that is a special case of the approach presented in [27] which in turn is a special case of an even more general approach presented in [38].
@@
The inverse problem we consider herein is of unique continuation type. There are many types of methods for this type of problem. In the framework we consider the earliest works considered quasi-reversibility [6]. The stabilized method we consider for unique continuation was first proposed in [7, 8] and [9]. More recent works use residual minimization in dual norm [16, 18, 10]. The optimal error estimates for unique continuation with a trace in a finite-dimensional space was first considered for Dirichlet trace in [12] and for Neumann trace in [13]. The idea of combining unique continuation in finite-dimensional space with collective data was first proposed in [25, 5] using linear algebra methods for the compression and direct solution of the linear unique continuation problem. Low rank solvers for the solution of inverse problems have also been designed in [36] using proper orthogonal decomposition.
In recent years, significant advancements have been made in utilizing machine learning for solving PDEs [26, 40, 35]. One important aspect is how to suitably and efficiently represent the learned solution [22, 39, 33, 3]. An application that comes very natural in the context of neural networks is the derivation of reduced order models [31, 37].
These developments are very useful in the context of inverse problems, where they have been utilized in both data- and model-driven inverse problems. In [32] a combination of networks and traditional methods is considered to recover the diffusion coefficient in Poisson’s and Burgers’ equations. In general, the same is done in [14] with the traditional method being FEM and the equations being elliptic and parabolic. Yet more examples of applying deep learning to this type of problem are given by [15, 4]. Anyone interested in the application of deep learning for PDE-solving has undoubtedly encountered PINNs [34] which are also used for inverse problems. Works not involving deep learning but still relevant are [28, 17] where projection-based reduced order models for inverse problems are presented. Taking the step to also include machine learning, some of the authors from the previous works give an overview of this mix in [24]. Another overview of using machine learning for inverse problems is given by [2]. In [23], an approach to reduce the error introduced by using operator learning for inverse problems is studied. As a contrast, [30] instead uses machine learning to reduce the error introduced by approximate forward models. Focusing instead on the other side of the computational spectrum, i.e., speed, [21] presents a physics-based deep learning methodology with applications to optimal control. The work [19] presents a modular machine learning framework for solving inverse problems in a latent space. Although using different techniques and approaches, this general description also holds for what we present here.
@@
In Section 2, we introduce the model problem and the finite element discretization; in Section 3, we present and prove stability and error estimates for a linear model problem; in Section 4, we develop a machine learning-based approach for solving the inverse problem; in Section 5, we present several numerical examples illustrating the performance of the method for various complexity of the given set of boundary data; and in Section 6 we summarize our findings and discuss future research directions.
@@ 2.1 Inverse Problem
$\boxed{\inf_{v\in V}\frac{1}{2}\|u_{0}-v\|_{L^{2}(\omega)}^{2}\quad\text{subject to}\quad\mathcal{P}(v)=0\quad\text{in $\Omega$}}$ (2.1)
@@
of observed Dirichlet data at the boundary. The dataset $\mathcal{G}$ may have different properties, but here we will assume that it is of the form
@@
$\boxed{\inf_{v\in V}\frac{1}{2}\|u_{0}-v\|_{L^{2}(\omega)}^{2}\quad\text{subject to}\quad\mathcal{P}(v)=0\quad\text{in $\Omega$},\quad v|_{\partial\Omega}\in\mathcal{G}}$ (2.5)
@@ 2.2 Finite Element Method
for $10\leq m\leq k\leq p+$ 1. Here $\lesssim$ means that there is a positive constant in the inequality (typically on the right-hand side) and $N(T)$ is the union of all elements that share a node with $T$.
@@
$\boxed{\inf_{v\in V_{h}}\frac{1}{2}\|u_{0}-v\|_{L^{2}(\omega)}^{2}\quad\text{subject to}\quad(\mathcal{P}(v),w)_{H^{-1}(\Omega),H^{1}(\Omega)}=0\quad\forall w\in V_{h,0},\quad v|_{\partial\Omega}\in\pi_{h}\mathcal{G}}$ (2.7)
@@ 3.1 The Continuous Problem
$g\in\mathcal{G}=\Bigl{\{}\sum_{n=1}^{N}a_{n}g_{n}\,|\,a_{n}\in\mathbb{R}\Bigr{\}}$ (3.2)
@@
$\mathcal{P}\varphi_{n}=0\quad\text{in $\Omega$},\qquad\varphi_{n}=g_{n}\quad\text{on $\partial\Omega$}$ (3.3)
@@
By unique continuation for harmonic functions, we conclude that $u_{1}-u_{2}$ is zero on the boundary and therefore $u_{1}=u_{2}$ since the set $\{g_{n}\}_{n=1}^{N}$ is linearly independent on $\partial\Omega$. It follows that $\{\varphi_{n}\}_{n=1}^{N}$ is linearly independent on $\omega$ and by finite dimensionality, there is a constant (it is $\lambda_{min}^{-1/2}$), such that
@@ 3.2 The Discrete Problem
The estimate (3.15) follows directly using the Cauchy–Schwarz inequality and the error estimates (3.12) for the approximate basis
@@ 3.3 Error Estimates
Let $m_{h}$ be defined by (3.23). Then, there is a constant, depending on $N$ but not $h$, such that
@@
    where we finally used the identity $I_{h}\widehat{v}=v$, which holds since $v\in\mathcal{U}_{N,h}$. Next, we bound the second term using the stabilizing terms in $m_{h}$. To that end, we observe that we have the orthogonality
@@
    we obtain using Galerkin orthogonality followed by partial integration
@@
    $\|e\|_{\omega}\lesssim h\|u_{0}\|_{\omega}$ (3.56)
@@
Observe that the stabilization is never explicitly used in order to obtain error estimates. Indeed its only role is to ensure the bound $\|\widehat{u}_{N,h}\|_{\mathbb{R}^{N}}\lesssim\|u_{N,h}\|_{m_{h}}$ without condition on the mesh.
Example (Exponential growth of the stability constant).
@@ 4 Methods Based on Machine Learning
  • Construct a parametrization of the data set by first approximately expanding the samples in a finite series of functions, for instance, using Proper Orthogonal Decomposition, and secondly using an autoencoder to find a possible nonlinear low-dimensional structure in the expansion coefficients.
@@
  • Composing the decoder, which maps the latent space to expansion coefficients, with the solution network, we obtain a differentiable mapping that can be used to solve the inverse problem in a lower-dimensional space.
@@ 4.1 Processing the Boundary Data
where $\widehat{g}_{n}=(g,\varphi_{n})_{\mathbb{R}^{N}}$. We introduce the mapping
@@
Letting $I_{N}$ denote the identity operator on $\mathbb{R}^{N}$, we have
@@
Next, we seek to find a possible nonlinear low-dimensional structure in the POD coefficients using an autoencoder $\phi_{\text{de}}\circ\phi_{\text{en}}$
@@
where $\phi_{\text{en}}$ denotes the encoder map and $\phi_{\text{de}}$ the decoder map. Letting $\mathbb{E}$ denote the expectation operator and $P$ an arbitrary probability distribution, the autoencoder is trained to minimize the loss
$\mathbb{E}_{\widehat{g}\sim P}\left[\|\widehat{g}-(\phi_{\text{de}}\circ\phi_{\text{en}})(\widehat{g})\|^{2}_{\mathbb{R}^{N}}\right]$ (4.6)
See Figure 2(a) for a schematic illustration. Here $Z\sim\mathbb{R}^{n_{Z}}$ is the latent space with dimension $n_{Z}<N$. If there is a low-dimensional structure, we may often take $n_{Z}$ significantly lower than $N$.
@@ 4.2 Operator Learning
The operator learning approach taken here is the same as in [27] which is a special case of a more general method presented in [38]. We discretize the PDE problem using finite elements and train a network
@@
which approximates the finite element solution to
$\mathcal{P}(u)=0\quad\text{in $\Omega$},\qquad u=\phi_{\mathrm{POD},N}^{\dagger}(\widehat{g})\quad\text{on $\partial\Omega$}$ (4.8)
see Figure 2(b). The output of the network is the finite element degrees of freedom (DoFs). For the training of the network we use the energy functional $E$ corresponding to the differential operator $\mathcal{P}$ as the foundation for the loss function. Again, letting $\mathbb{E}$ denote the expectation operator and $P$ an arbitrary probability distribution, the loss function that we minimize during training is
$\mathbb{E}_{\widehat{g}\sim P}\left[E(\phi_{u,N,h}(\widehat{g}))\right]$ (4.9)
If there is no corresponding energy functional, one can instead minimize the residual of the finite element problem. It should be noted though, that assembling the residual instead of the energy has a greater computational cost and that the residual is not as easily and naturally decomposed into its local contributions as the energy. For technical details about network architecture and training used in this work, we refer to Section 5.2.